1/3

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

Der Werwolf The Werewolf The Banshee (An Approach)

El Hechice- ro

translated by Alexander Gross translated by Max Knight translated by Hans Eiseneck

Ein Werwolf eines

Nachts entwich von Weib und Kind

und sich begab an eines Dorfschulleh-

rers Grab und bat ihn: �Bitte,

beuge mich!“ A Werewolf, troubled by

his name, Left wife and brood one

night and came To a hidden graveyard to

enlist The aid of a long-dead phi-

lologist.

One night, a banshee

slunk away from mate and child,

and in the gloom went to a village tea-

cher’s tomb, requesting him: “Inflect

me, pray.”

Un hechicero, con

esposa vino un día a la

fosa de su maestro, y

al profesor pidió: “¿Me cuen-

tas? ¡Por favor!”

Der Dorfschulmeister

stieg hinauf auf seines Blech-

schilds Messingknauf

und sprach zum Wolf,

der seine Pfoten geduldig kreuzte vor

dem Toten:

“Oh sage, wake up, please

don’t berate me,” He howled sadly, “Just con-

jugate me.” The seer arose a bit unstea-

dy Yawned twice, wheezed on-

ce, and then was ready.

The village teacher

climbed up straight upon his grave stone

with its plate and to the apparition

said who meekly knelt befo-

re the dead:

El profesor sin va-

cilar subió y comenzó a

hablar. Desde allá de su

letrero se dirigió al hechi-

cero:

�Der Werwolf“ – sprach der gute Mann,

�des Werwolfs, Geni-

tiv sodann, dem Werwolf, Dativ,

wie man’s nennt, den Werwolf, - damit

hat’s ein End!“ “Well, ‘Werewolf’ is your

plural past, While ‘Waswolf’ is singu-

larly cast: There’s ‘Amwolf’ too, the

present tense, And ‘Iswolf,’ ‘Arewolf’ in

this same sense.”

“The banSHEE, in the

subject’s place; the banHERS, the pos-

sessive case. The banHER, next, is

what they call objective case–and that

is all.”

“El hechiuno” le

explica, “el hechidos” de-

spués indica, “el hechitres” dice

al fin, “con eso basta

¡chiquitín!”

Dem Werwolf schmei-

chelten die Fälle er rollte seine Au-

genbälle. �Indessen“, bat er,

�füge doch zur Einzahl auch die

Mehrzahl noch!“ “I know that–I’m no mental

cripple– The future form and parti-

ciple Are what I crave,” the be-

ast replied. The scholar paused–again

he tried:

The banshee marveled

at the cases and writhed with plea-

sure, making faces, but said: “You did not

add, so far, the plural to the singu-

lar!”

El hechicero, con-

movido por lo que capta

su oído, al maestro ruega,

encantado, que le adjunte un

quebrado.

…  …  …  …

2/3

Der Dorfschullehrer

aber musste gestehn, dass er von

ihr nichts wusste. Zwar Wölfe gab’s in

großer Schar, doch �Wer“ gab’s nur

im Sinular.

“A ‘Will-be-wolf?’ It’s just

too long: ‘Shall-be-wolf?’ ‘Has-been-

wolf?’ Utterly wrong! Such words are wounds

beyond all suture– I’m sorry, but you have no

future.”

The teacher, though,

admitted then that this was not wi-

thin his ken. “While bans are fre-

quent”, he advised, “A she cannot be pluri-

zed.”

Pero el profesor

confiesa que nada sabe con

certeza: “Hay muchos he-

chos. Hechiceros

únicamente hay enteros.”

Der Wolf erhob sich

tränenblind – er hatte ja doch Weib

und Kind! Doch da er keine Ge-

lehrter eben, so schied er dankend

und ergeben.

The Werewolf knew better–his sons still slept

At home, and homewards

now he crept, Happy, humble, without

apology For such folly of philology.

The banshee, rising

clammily, wailed: “What about

my family?” Then, being not a lear-

ned creature, said humbly “Thanks”

and left the teacher.

El brujo se volvió

muy triste: “Mi cara mitad,

¿no la viste?” Ya que no era sa-

bio con gratitud se

despidió.

…  …  …  …

3/3

Author of this German poem is Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_ url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into hispana by Hans Eiseneck.

Arg-825-1656 (2013-02-13 18:46:00)

Tiu ĉi trsduko en la kastilian lingvon troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .